OpenStudy (anonymous):

Rule of correspondance for relations I do know how to do it (mostly) but I am having trouble with a few. I guess the question is can I use negative numbers to solve them? {(x,y): (1,2),(2,5), (3,8), , ,}

OpenStudy (mathstudent55):

Can you tell if the relation is linear?

OpenStudy (anonymous):

You mean going in a straight line on a graph?

OpenStudy (mathstudent55):

Yes. Can you tell if its graph is a straight line?

OpenStudy (anonymous):

my graph looks awful but yes it seems to be straight or linear.

OpenStudy (mathstudent55):

Ok. I'll tell you how you can tell if it is a line without graphing. A linear relation has the following. For an equal difference in x, there must equal difference in y.

OpenStudy (mathstudent55):

I'll explain how that is used.

OpenStudy (mathstudent55):

You have only 3 points given. (1, 2), (2, 5), (3, 8)

OpenStudy (mathstudent55):

Look at the first two points: (1, 2) and (2, 5) The difference in x-coordinates is 2 - 1 = 1 The difference in y-coordinates is 5 - 2 = 3 Ok so far?

OpenStudy (anonymous):

Well I also noticed the pattern it had. one right, three up. if that's what you mean.

OpenStudy (mathstudent55):

Exactly. That shows the same slope between points which means a line. Great observation.

OpenStudy (anonymous):

:D

OpenStudy (mathstudent55):

The difference in x was 1 and the difference in y was 3. Now let's look at the last two points. (2, 5), (3, 8) The difference is x is 3 - 2 = 1 The difference in y is 8 - 5 = 3 Again we have a difference in x of 1 and a difference in y of 3. Since for the same difference in x (1), we get teh same difference in y(3), we are dealing with a straight line.

OpenStudy (mathstudent55):

Now that we know we have a straight line, we can find the equation of the line.

OpenStudy (mathstudent55):

Do you know how to find the equation of a line given two points?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Not really

OpenStudy (mathstudent55):

Ok, no problem. I'll explain it to you. There are several methods, but here is one of them.

OpenStudy (mathstudent55):

The two-point equation of a line A line that passes through points \((x_1, y_1)\) and \((x_2, y_2)\) has equation \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}(x - x_1)\)

OpenStudy (mathstudent55):

You can use any two points of your three given points. Let's use the first two points (1, 2) and (2, 5)

OpenStudy (anonymous):

sorry my tracad is messing around. but the answer appears to be 3 over 1?

OpenStudy (mathstudent55):

\(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}(x - x_1)\) \(y - 2 = \dfrac{5 - 2}{2 - 1}(x - 1)\) \(y - 2 = \dfrac{3}{1}(x - 1)\) \(y - 2 = 3(x - 1)\) \(y - 2 = 3x - 3)\) \(y = 3x - 1\) The equation gives you the rule of the relation. Take the x value, multiply by 3, and subtract 1. Try it with your 3 points, and you'll see that it works.

OpenStudy (anonymous):

Ok I see what you have there. If I try one with a different equation can you check to see I'm doing it right? Also can this only be done with linear equations?

OpenStudy (anonymous):

@mathstudent55 (2,3) (4,4) (6,5) y=x-1? or did I miss something?

OpenStudy (anonymous):

Actually that can't be right...

OpenStudy (mathstudent55):

\(y = \dfrac{1}{2}x + 2\)

OpenStudy (mathstudent55):

Sorry, but gtg.

OpenStudy (anonymous):

ok thanks though!

OpenStudy (mathstudent55):

yw